# A Unified Pulse-Shaped OFDM Framework for Chirp-Domain Waveforms: Continuous-Time Modeling and Practical I/O Analysis

### 15 Mar 2026 v1

@@ Abstract
In this paper, a unified framework for chirp-domain waveforms, including orthogonal chirp division multiplexing (OCDM) and affine frequency division multiplexing (AFDM), is developed. Based on their continuous-time representations, we show that these waveforms fall within the conventional Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where the root chirp corresponds directly to the prototype pulse in the WH framework. Since the chirp is a constant-envelope signal and is transparent to subcarrier orthogonality, these waveforms can be further interpreted as pulse-shaped (PS) orthogonal frequency division multiplexing (OFDM). Within the developed PS-OFDM framework, the power spectral density of chirp-domain waveforms is derived analytically. We then discuss existing practical implementations of chirp-domain waveforms, which rely on sub-Nyquist discrete-time samples and therefore exhibit frequency aliasing. The resulting aliased waveform is analyzed, and the orthogonality among the embedded aliased chirps is discussed. It is shown that the aliased chirps are conditionally orthogonal, whereas the implemented approximate aliased chirps can maintain mutual orthogonality when an appropriate sample-wise pulse-shaping filter is applied. We further derive an exact input-output relation for the implemented chirp-domain waveform over a delay-Doppler (DD) channel, showing that the effective channel observed at a practical receiver does not, in general, admit a DD spreading-function model commonly assumed in the literature. The implementation complexity is also investigated and compared with that of orthogonal delay-Doppler division multiplexing (ODDM), the DD-domain MC waveform defined within the evolved WH framework. Finally, simulation results are provided to verify the analysis.
@@ I Introduction
Next-generation wireless systems, including sixth-generation (6G) mobile networks, vehicle-to-everything (V2X) communications, high-speed railway links, and millimeter-wave/terahertz (mmWave/THz) systems, are increasingly expected to operate in scenarios with very high user or scatterer mobility. Such mobility induces a large Doppler spread and shortens the channel coherence time. In conventional orthogonal frequency-division multiplexing (OFDM) systems, this often manifests as rapidly increasing inter-carrier interference (ICI): once the normalized Doppler spread (e.g., the maximum Doppler shift divided by the subcarrier spacing) is no longer sufficiently small, subcarriers may effectively lose their frequency-domain orthogonality. This behavior is consistent with the fact that OFDM is typically designed under an approximately linear time-invariant (LTI) channel assumption, whereas under high mobility the physical channel is more appropriately modeled as a linear time-varying (LTV) system. Consequently, it becomes necessary to represent the channel in domains that better capture the LTV structure, and to design waveforms and transceivers accordingly.
@@
Since the physical units of delay and Doppler are time and frequency, respectively, the DD domain is naturally a time-frequency (TF) domain. As a result, it has been shown that DD-domain waveforms can be interpreted within the Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where ODDM evolves the existing WH framework by introducing a (bi)orthogonal WH-subset-based design [8, 9]. Recent studies of ODDM waveforms over general physical channels can also be found in [10, 11, 12].
On the other hand, although orthogonal chirps play a central role in chirp-based waveforms, a systematic framework for understanding the fundamentals of these orthogonal-chirp-based waveforms is still lacking. Moreover, existing studies of chirp-domain waveforms are often based on discrete-time sequences generated by the discrete Fresnel transform (DFnT) [6] or the discrete affine Fourier transform (DAFT)[7], whereas the transmitted waveform is ultimately a continuous-time signal that propagates through a general physical channel with continuous-valued delay and Doppler. This discrepancy may lead to a mismatch between discrete-sequence-level analyses and the underlying channel behavior characterized by continuous-valued delay and Doppler.
In this paper, we develop a unified framework for chirp-domain waveforms, including OCDM and AFDM. We show that these waveforms also fall within the conventional WH framework for MC waveforms, where the root chirp corresponds directly to the prototype pulse in the WH framework. Under this interpretation, chirp-domain waveforms can be viewed as pulse-shaped (PS) OFDM with a constant-envelope prototype pulse. We also analyze practical implementations based on sub-Nyquist discrete-time samples, investigate the resulting aliased chirps and their orthogonality properties, and derive the exact input-output (I/O) relation over DD channels. The contributions of this paper are summarized as follows:
@@
  • For implemented waveforms containing approximate aliased chirps, we derive the exact I/O relation over DD channels. Our analysis shows that the effective channel observed at the receiver does not, in general, admit the conventional DD spreading-function representation commonly assumed in the literature. As a result, compared with existing studies that rely solely on discrete-sequence models and may therefore deviate from the behavior observed in practical transceivers, the derived I/O relation provides a more accurate characterization of the practical performance of chirp-domain waveforms.
  • Simulations are provided to verify the theoretical analysis, validate the derived PSD of chirp-domain waveforms, and confirm the derived I/O relation. We also analyze the implementation complexity of chirp-domain waveforms and compare it with that of ODDM. When occupying the same TF resources, the implementation complexity of ODDM is shown to be significantly lower than that of chirp-domain waveforms.
The remainder of this paper is organized as follows. In Section II, beginning with a general representation of ditigal modulation waveforms, the WH framework for MC waveforms is introduced. Section III presents OCDM and AFDM in their continuous-time forms. Based on this, the unified PS-OFDM framework for chirp-domain waveforms is developed and the PSD is derived. In Section IV, practical implementations based on sub-Nyquist samples are investigated; the ideal and approximate aliased chirps are presented, and their orthogonality properties are studied. Section V derives the I/O relation of implemented chirp-domain waveforms with approximate aliased chirps over DD channels and verifies the effective channel observed at the receiver. Simulation validation and complexity comparison are also provided in Section VI. Finally, Section VII concludes the paper.
Notations: In this paper, uppercase boldface letters are used to represent matrices, and lowercase boldface letters are used for column vectors. The superscript $\mathsf{T}$, $\mathsf{H}$ and * denote the transpose, conjugate transpose, and conjugate operators, respectively. Also, $\Pi_{T}(t)$ denotes the rectangular function of unit energy and support of $[0,T)$, while $\mathcal{A}_{g,\tilde{g}}(\tau,\nu)$ stands for the cross ambiguity function between two pulses $g(t)$ and $\tilde{g}(t)$, given by
$\mathcal{A}_{g,{\tilde{g}}}(\tau,\nu)$ $=\langle g(t),{\tilde{g}}(t-\tau)e^{j2\pi\nu(t-\tau)}\rangle,$
@@ II MC Waveform Principles
Given the symbol index $i\in\mathbb{Z}$, a real-valued passband digital modulation waveform with carrier frequency $f_{c}$ and phase $\theta_{c}$ is typically represented as [13, 14, 15, 9]:
$x_{pb}(t)=\sum_{i}\sqrt{2}A_{i}\cos(2\pi(f_{c}+f_{i}(t))t+\theta_{c}+\theta_{i})\rho_{i}(t),$
@@
To avoid inter-symbol interference (ISI), the transmit pulses $\{g_{i}(t),i\in\mathbb{Z}\}$ are required to be mutually orthogonal. Hence, a digital modulation waveform is defined by a set of orthogonal pulses, also referred to as basis functions.
For MC systems, the symbol index $i$ is replaced by a time index $m$ and a frequency index $n$, and the MC waveform is given by [8, 9, 16]:
@@
For such an MC waveform characterized by a TF grid $(\Delta T,\Delta F)$ and a prototype pulse $g(t)$, a design framework based on the WH frame theory is well-known[16], as $g_{m,n}(t)$ in (3) is the WH or Gabor function adopted in TF signal analysis [17].
@@ III-A OCDM Waveform
for $0\leq k\leq N-1$, where $e^{\frac{j\pi}{4}}e^{-j\frac{\pi}{N}(k-n)^{2}}$ is exactly the $(k,n)$-th entry of the $N\times N$ IDFnT matrix up to a scaling factor.
@@ III-B AFDM Waveform
By sampling (11) over both $t$ and $u$ with intervals $\Delta t$ and $\Delta u$, respectively, the discrete AFT (DAFT) and the corresponding inverse DAFT can be defined under the condition [20]
@@
where $N$ denots denotes the number of samples over $u$. For the commonly adopted symmetric sampling grid with $\Delta t=\Delta u$, the transform parameter $b$ follows
@@
In [7], the inverse DAFT is used to generate the discrete-time AFDM sequence as
@@
Bearing in mind that $\Delta t=\frac{T}{N}$, and comparing (20) with (7), it is clear that OCDM is a special case of AFDM with $c_{1}=\frac{1}{2N}.$ Thus, we hereafter use
@@ III-C WH and PS-OFDM Frameworks
Recall that when $g(t)$ forms an orthogonal WH set over the TF grid $\left(\Delta T,\Delta F\right)$, the MC waveform in (3) becomes a PS-OFDM signal. As a result, chirp-domain waveforms not only belong to the WH framework but can also be interpreted as PS-OFDM signals with a constant-envelope prototype pulse.
@@ III-D1 PSD of PS-OFDM Signals
Assume that the information-bearing symbols $X[m,n]$ are i.i.d. with zero mean and average power $\sigma_{X}^{2}=\mathbb{E}[|X[m,n]|^{2}]$. It can be proved that $x(t)$ is wide-sense cyclostationary with period $\Delta T$. Given the time-average autocorrelation function of cyclostationary signal is defined as [14] :
@@ III-D2 PSD of AFDM Waveform
$S_{\rm AF}(f)=\frac{\sigma_{X}^{2}}{NT}\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}\left|G_{\rm AF}\left(f-\frac{n}{T}\right)\right|^{2},$ (31)
where $e^{j2\pi c_{2}n^{2}}$ is a phase factor that does not affect the second-order statistics of $\dot{X}[n]$, and $G(f)$ is the corresponding prototype pulse spectrum given by
@@ IV-B Approximate Aliased AFDM Waveform
A discrete sequence has infinite bandwidth, and therefore must be pulse-shaped into a analog waveform that is band-limited before being carrier-modulated and transmitted over the air.
@@
In the context of MC waveforms, pulse shaping can be somewhat confusing because an MC symbol usually spans multiple time-domain samples, whereas in single-carrier modulation one baud-rate sample corresponds to one symbol. Pulse shaping in MC waveforms typically involves two steps: first, interpolating discrete-time samples using a filter (equivalent to the sample-wise pulse shaping used in the single-carrier case), and then multiplying the result by the symbol-wise prototype pulse (symbol-wise pulse shaping). When the prototype pulse is constant-envelope, such as the rectangular pulse, sample-wise pulse shaping alone may be sufficient, subject to appropriate parameter settings [8, 9].
@@
for $0\leq k\leq N-1$. The chirp sequence can be represented as a vector $\bm{\phi}_{n}=[\phi_{n}[0],\cdots,\phi_{n}[N-1]]^{T},$ corresponding to the $n$-th column of the IDAFT matrix. Since $\langle\bm{\phi}_{n},\bm{\phi}_{n^{\prime}}\rangle=\delta(n-n^{\prime})$, it is well known that
$\langle\phi_{n}^{(a)}(t),\phi_{n^{\prime}}^{(a)}(t)\rangle=\delta(n-n^{\prime}),$
when $a(t)$ is a root-Nyquist pulse for the Nyquist interval $\frac{T}{N}$ [15].
@@ V I/O Relation over DD channels
Assume the LTV channel consists of $P$ propagation paths. For the $p$-th path, let $h_{p}$, $\tau_{p}$, and $\nu_{p}$ denote its complex gain, delay, and Doppler shift, respectively. This channel model is considered valid during the channel’s ”stationarity” interval[1, 3, 8]. Its representation in the DD domain, also known as the DD spreading function, is given by
@@
and the delay spread of the channel is given by $\mathcal{T}=\tau_{p}-\tau_{1}<T$. For a transmitted baseband signal $x(t)$, the received baseband signal $r(t)$ is expressed as
@@ V-A Discrete-Sequence Based I/O Relation
The I/O relation of AFDM waveform has been derived in [7], where each path is assumed with an on-grid delay that can be normalized to an integer as $l_{p}=\tau_{p}\frac{N}{T}.$ Then, the received samples are given as [7]
@@
$h_{k^{\prime},l}$ $=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}k^{\prime}}\delta(l-l_{p})$ (43)
is the time-varying impulse response of channel at time $k^{\prime}$ and delay $l$. A chirp-periodic prefix (CPP) with length $L_{\rm CPP}>\lceil{\mathcal{T}}\frac{N}{T}\rceil-1$ is ususally prepended into $x_{AF}[k]$ in (42).
After discarding the CPP, (42) can be written in the matrix form
@@
with $\mathbf{z}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$,
$\mathbf{H}=\sum_{p=1}^{P}h_{p}\mathbf{\Gamma}_{p}\mathbf{\Delta}_{\nu_{p}}\mathbf{\Pi}^{l_{p}},$ (45)
and $\mathbf{\Pi}$ is the $N\times N$ forward cyclic-shift matrix, $\mathbf{\Delta}_{\nu_{p}}\triangleq\mathrm{diag}\left(1,e^{j2\pi\nu_{p}},\cdots,e^{j2\pi(N-1)\nu_{p}}\right)$ and $\mathbf{\Gamma}_{p}$ is a $N\times N$ diagonal matrix given by
@@
By applying the DAFT matrix $\mathbf{A}$ to $\mathbf{y}$, the DAFT domain I/O relation becomes
@@
The DAFT domain I/O matrix $\mathbf{H}_{u}$ plays a fundamental role in performance evaluation and AFDM waveform design, such as in selecting the parameters $c_{1}$ and $c_{2}$ to achieve optimal diversity order and reduced pilot overhead [7].

### 31 Mar 2026 v2

@@ Abstract
In this paper, a unified framework for chirp-domain waveforms, including orthogonal chirp division multiplexing (OCDM) and affine frequency division multiplexing (AFDM), is developed. Based on their continuous-time representations, we show that these waveforms fall within the conventional Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where the root chirp corresponds to the prototype pulse in the WH framework. Since the chirp is a constant-envelope signal and is transparent to subcarrier orthogonality, these waveforms can be further interpreted as pulse-shaped (PS) orthogonal frequency division multiplexing (OFDM). Within the developed PS-OFDM framework, the power spectral density of chirp-domain waveforms is derived analytically. We then discuss existing practical implementations of chirp-domain waveforms, which rely on sub-Nyquist discrete-time samples and therefore exhibit frequency aliasing. The resulting aliased waveform is analyzed, and the orthogonality among the embedded aliased chirps is discussed. It is shown that the aliased chirps are conditionally orthogonal, whereas the implemented approximate aliased chirps can maintain mutual orthogonality when an appropriate sample-wise pulse-shaping filter is applied. We further derive an exact input-output (I/O) relation for the implemented chirp-domain waveform over delay-Doppler (DD) channels, showing that the effective channel at a practical receiver does not, in general, conform to a superposition of pure path-wise DD components, resulting in a non-negligible deviation from the I/O relation commonly used in the literature. The implementation complexity is also investigated and compared with that of orthogonal delay-Doppler division multiplexing (ODDM), the DD-domain MC waveform defined within the evolved WH framework. Finally, simulation results are provided to verify the analysis.
@@ I Introduction
Next-generation wireless systems, including sixth-generation (6G) mobile networks, vehicle-to-everything (V2X) communications, high-speed railway links, and millimeter-wave/terahertz (mmWave/THz) systems, are increasingly expected to operate in scenarios with very high user or scatterer mobility. Such mobility induces a large Doppler spread and shortens the channel coherence time. In conventional orthogonal frequency division multiplexing (OFDM) systems, this often manifests as rapidly increasing inter-carrier interference (ICI). Once the normalized Doppler spread (e.g., the maximum Doppler shift divided by the subcarrier spacing) is no longer sufficiently small, subcarriers may effectively lose their frequency-domain orthogonality. This behavior is consistent with the fact that OFDM is typically designed under an approximately linear time-invariant (LTI) channel assumption, whereas under high mobility the physical channel is more appropriately modeled as a linear time-varying (LTV) system. Consequently, it becomes necessary to represent the channel in domains that better capture the LTV structure, and to design waveforms and transceivers accordingly.
@@
Since the physical units of delay and Doppler are time and frequency, respectively, the DD domain is naturally a time-frequency (TF) domain. As a result, it has been shown that DD-domain waveforms can be interpreted within the well-known Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where ODDM evolves the existing WH framework by introducing a (bi)orthogonal WH-subset-based design [8, 9, 10]. Recent studies of ODDM waveforms over general physical channels can be found in [11, 12, 13].
On the other hand, although orthogonal chirps play a central role in chirp-based waveforms, a systematic framework for understanding their fundamental properties remains elusive. Furthermore, while some studies have addressed the continuous-time representation of chirp-domain waveforms [6, 14, 15], much of the existing literature is based on discrete-time sequences generated by the discrete Fresnel transform (DFnT) [6] or the discrete affine Fourier transform (DAFT) [7]. Given that the transmitted waveform is ultimately a continuous-time signal propagating through a physical channel characterized by continuous-valued delay and Doppler, this discrepancy can lead to a non-negligible mismatch between discrete-sequence-level analyses and actual channel behavior.
In this paper, we develop a unified framework for chirp-domain waveforms, including OCDM and AFDM. We show that these waveforms also fall within the conventional WH framework for MC waveforms, where the root chirp corresponds to the prototype pulse in the WH framework. Under this interpretation, chirp-domain waveforms can be viewed as pulse-shaped (PS) OFDM with a constant-envelope prototype pulse. We also analyze practical implementations based on sub-Nyquist discrete-time samples, investigate the resulting aliased chirps and their orthogonality properties, and derive the exact input-output (I/O) relation over DD channels. The contributions of this paper are summarized as follows:
@@
  • For the implemented waveform containing approximate aliased chirps, we derive the exact I/O relation over DD channels. Our analysis shows that the effective channel at a practical receiver does not, in general, conform to a superposition of pure path-wise DD components commonly assumed in the literature. As a result, compared with existing studies that rely solely on discrete-sequence models and may therefore deviate from the behavior observed in practical transceivers, the derived I/O relation provides a more accurate characterization of the practical performance of chirp-domain waveforms.
  • Simulations are provided to verify the theoretical analysis, validate the derived PSD, and confirm the derived I/O relation. We also analyze the implementation complexity of chirp-domain waveforms and compare it with that of ODDM. We show that when transmitting the same number of symbols with the same subcarrier spacing, ODDM exhibits notably lower implementation complexity than chirp-domain waveforms.
The remainder of this paper is organized as follows. In Section II, beginning with a general representation of digital modulation waveforms, the WH framework for MC waveforms is introduced. Section III presents OCDM and AFDM in their continuous-time forms. Based on this, the unified PS-OFDM framework for chirp-domain waveforms is developed and the PSD is derived. In Section IV, practical implementations based on sub-Nyquist samples are investigated; the ideal and approximate aliased chirps are presented, and their orthogonality properties are studied. Section V derives the I/O relation of the implemented chirp-domain waveform with approximate aliased chirps over DD channels and identifies the effective channel observed at a practical receiver. Simulation validation and complexity comparison are provided in Section VI. Finally, Section VII concludes the paper.
Notations: Uppercase and lowercase boldface letters represent matrices and column vectors, respectively. The superscripts $\mathsf{T}$, $\mathsf{H}$, and * indicate transpose, conjugate transpose, and complex conjugate. The operator $\star$ denotes convolution. The function $\Pi_{T}(t)$ is a unit-energy rectangular pulse with support $[0,T]$, and $\mathcal{A}_{g,\tilde{g}}(\tau,\nu)$ denotes the cross-ambiguity function between $g(t)$ and $\tilde{g}(t)$, given by
$\mathcal{A}_{g,\tilde{g}}(\tau,\nu)$ $=\langle g(t),\tilde{g}(t-\tau)e^{j2\pi\nu(t-\tau)}\rangle$
@@ II MC Waveform Principles
Given the symbol index $i\in\mathbb{Z}$, a real-valued passband digital modulation waveform with carrier frequency $f_{c}$ and phase $\theta_{c}$ is typically represented as [16, 17, 18]:
$x_{pb}(t)=\sqrt{2}\sum_{i}A_{i}\cos(2\pi(f_{c}+f_{i}(t))t+\theta_{c}+\theta_{i})\rho_{i}(t),$
@@
To avoid inter-symbol interference (ISI), the transmit pulses $\{g_{i}(t),i\in\mathbb{Z}\}$ are required to be mutually orthogonal. Hence, a digital modulation waveform can be defined by a set of orthogonal pulses, also referred to as basis functions.
For MC systems, the symbol index $i$ is replaced by a time index $m$ and a frequency index $n$, and the MC waveform is given by [9, 10, 19, 20]:
@@
For such an MC waveform characterized by a TF grid $(\Delta T,\Delta F)$ and a prototype pulse $g(t)$, a design framework based on the WH frame theory is well-known[19, 20], as $g_{m,n}(t)$ in (3) is the WH or Gabor function adopted in TF signal analysis [21].
@@ III-A OCDM Waveform
for $0\leq k\leq N-1$, where $e^{\frac{j\pi}{4}}e^{-j\frac{\pi}{N}(k-n)^{2}}$ is exactly the $(k,n)$-th entry of the $N\times N$ inverse DFnT (IDFnT) matrix up to a scaling factor.
@@ III-B AFDM Waveform
By sampling (11) over both $t$ and $u$ with intervals $\Delta t$ and $\Delta u$, respectively, the discrete AFT (DAFT) and the corresponding inverse DAFT (IDAFT) can be defined under the condition [24]
@@
where $N$ denotes the number of samples over $u$. For the commonly adopted symmetric sampling grid with $\Delta t=\Delta u$, the transform parameter $b$ follows
@@
In [7], the IDAFT is used to generate the discrete-time AFDM sequence as
@@
Bearing in mind that $\Delta t=\frac{T}{N}$, and comparing (20) with (7), it is clear that OCDM is a special case of AFDM with $c_{1}=-\frac{1}{2N}.$ Thus, we hereafter use
@@ III-C WH and PS-OFDM Frameworks
Recall that when $g(t)$ is non-rectangular and forms an orthogonal WH set over the TF grid $\left(\Delta T,\Delta F\right)$, the MC waveform in (3) becomes a PS-OFDM signal. Consequently, chirp-domain waveforms not only fall within the WH framework but can also be interpreted as PS-OFDM signals with a constant-envelope prototype pulse.
@@ III-D1 PSD of PS-OFDM Signals
Assume that the information-bearing symbols $X[m,n]$ are i.i.d. with zero mean and average power $\sigma_{X}^{2}=\mathbb{E}[|X[m,n]|^{2}]$. It can be proved that $x(t)$ is wide-sense cyclostationary with period $\Delta T$. Given the time-average autocorrelation function of cyclostationary signal defined as [17] :
@@ III-D2 PSD of AFDM Waveform
$S_{\rm AF}(f)=\frac{\sigma_{X}^{2}}{NT}\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}\left|G_{\rm AF}\left(f-\frac{n+N/2}{T}\right)\right|^{2},$ (31)
where $e^{j2\pi c_{2}n^{2}}$ in (22) is a phase factor that does not affect the second-order statistics of $\dot{X}[n]$, and $G(f)$ is the corresponding prototype pulse spectrum given by
@@ IV-B Approximate Aliased AFDM Waveform
A discrete-time sequence has infinite bandwidth and must therefore be pulse-shaped into a band-limited analog waveform before carrier modulation and over-the-air transmission. It should be noted that generating a discrete-time sequence followed by pulse shaping is only one possible implementation method. The modulation waveform can also be generated directly without explicitly forming the intermediate sequence.
@@
In MC waveforms, pulse shaping is more complex because each symbol spans multiple time-domain samples usually generated by the inverse fast Fourier transform (IFFT), unlike single-carrier modulation where each sample corresponds to one symbol. It typically involves two steps: sample-wise pulse shaping via interpolation filtering (equivalent to that used in the single-carrier case), and subsequent symbol-wise pulse shaping via multiplication with a prototype pulse. When the prototype pulse has a constant envelope (e.g., rectangular), sample-wise pulse shaping alone may suffice with appropriate parameter settings [9, 10]. Such waveforms are also referred to as filtered OFDM [25].
@@
for $0\leq k\leq N-1$. The chirp sequence can be represented as a vector $\bm{\phi}_{n}=[\phi_{n}[0],\cdots,\phi_{n}[N-1]]^{\mathsf{T}},$ corresponding to the $n$-th column of the IDAFT matrix. Since $\langle\bm{\phi}_{n},\bm{\phi}_{n^{\prime}}\rangle=\delta(n-n^{\prime})$, it is well known that
$\langle\phi_{n}^{(a)}(t),\phi_{n^{\prime}}^{(a)}(t)\rangle=N\delta(n-n^{\prime}),$
when $a(t)$ is a root-Nyquist pulse with Nyquist interval $\frac{T}{N}$ [18].
@@ V I/O Relation over DD channels
Assume the LTV channel consists of $P$ propagation paths. For the $p$-th path, let $h_{p}$, $\tau_{p}$, and $\nu_{p}$ denote its complex gain, delay, and Doppler shift, respectively. This channel model is considered valid during the channel’s “stationary” interval[1, 3, 9]. Its representation in the DD domain, also known as the DD spreading function, is given by
@@
Then, the relative delay and the delay spread of the channel are given by $\tilde{\tau}_{p}=\tau_{p}-\tau_{1}$ and $\tilde{\tau}_{P}<T$, respectively. For a transmitted baseband signal $x(t)$, the received baseband signal $y(t)$ is usually expressed as
@@ V-A Discrete-Sequence Based I/O Relation
The I/O relation of AFDM waveform has been derived in [7], where each path is assumed with an on-grid delay that can be normalized to an integer as $l_{p}=\tau_{p}\frac{N}{T}.$ Then, the received samples are given as
@@
$h_{k^{\prime},l}$ $=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}k^{\prime}\frac{T}{N}}\delta(l-l_{p})$ (43)
is the time-varying impulse response of channel at time $k^{\prime}$ and delay $l$. $x_{AF}[\cdot]$ in (42) usually includes a chirp-periodic prefix (CPP) with length $L_{\rm CPP}>\lceil{\tilde{\tau}_{P}}\frac{N}{T}\rceil-1$.
In the receiver, after discarding the CPP and letting $\mathbf{y}=[y[0],\cdots,y[N-1]]^{\mathsf{T}}$, $\mathbf{x}=[x_{\rm AF}[0],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, (42) can be expressed in matrix form as
@@
where $\mathbf{z}=[z[0],\cdots,z[N-1]]^{\mathsf{T}}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$, and
$\mathbf{H}=\sum_{p=1}^{P}h_{p}\mathbf{\Gamma}_{p}\mathbf{\Delta}_{\nu_{p}}\mathbf{\Pi}^{l_{p}}.$ (45)
In this expression, $\mathbf{\Pi}$ denotes the $N\times N$ forward cyclic-shift matrix, $\mathbf{\Delta}_{\nu_{p}}\triangleq\mathrm{diag}\left(1,e^{j2\pi\nu_{p}},\cdots,e^{j2\pi(N-1)\nu_{p}}\right)$, and $\mathbf{\Gamma}_{p}$ is a $N\times N$ diagonal matrix given by
@@
By applying the $N\times N$ DAFT matrix $\mathbf{A}$ to $\mathbf{y}$, the DAFT domain I/O relation becomes
@@
The DAFT domain I/O matrix $\mathbf{H}_{u}$ plays a fundamental role in AFDM waveform design and performance evaluation, such as in selecting the parameters $c_{1}$ and $c_{2}$ to achieve optimal diversity order and reduced pilot overhead [7].